# COLT: Enhancing Video Large Language Models with Continual Tool Usage

### v2

@@ Abstract
The success of Large Language Models (LLMs) has significantly propelled the research of video understanding. To harvest the benefits of well-trained expert models (i.e., tool), video LLMs prioritize the exploration of tool usage capabilities. Existing methods either prompt closed-source LLMs or employ the instruction tuning paradigm for tool-use finetuning. These methods, however, assume an established repository of fixed tools and struggle to generalize to real-world environments where tool data is perpetually evolving and streaming in. To this end, we propose to enhance open-source video LLMs with COntinuaL Tool usage (termed COLT), which automatically acquires tool-use ability in a successive tool stream without suffering “catastrophic forgetting” of the past learned tools. Specifically, our COLT incorporates a learnable tool codebook as a tool-specific memory system. Then relevant tools are dynamically selected based on the similarity between user instruction and tool features within the codebook. To unleash the tool usage potential of video LLMs, we collect a video-centric tool-use instruction tuning dataset VideoToolBench. Extensive experiments on both previous video LLM benchmarks and the tool-use-specific VideoToolBench dataset demonstrate the state-of-the-art performance of our proposed COLT.
Introduction
Large Language Models (LLMs), e.g., GPT (Achiam et al. 2023; Brown et al. 2020), PaLM (Anil et al. 2023) and LLaMA (Touvron et al. 2023a, b), have exhibited remarkable success in understanding user instructions, aligning with human intents and generating trustworthy responses. Taking one step further, recent endeavors in multi-modal LLMs (Li et al. 2023a; Alayrac et al. 2022; Zhu et al. 2023b; Liu et al. 2024a; Zhang et al. 2023a; Wu et al. 2023; Li et al. 2024b; Zhan et al. 2024) augment LLMs with the capacity of multi-modal perception, which has aroused significant research interest within the academic community and beyond.
Figure 1: Our proposed COLT continually learns to invoke tools from tool-stream data without catastrophic forgetting. Benefiting from tool usage, COLT (a) is adept at dynamic content understanding and (b) supports flexible generation compared to existing methods (Lin et al. 2023a; Li et al. 2023c). The incorrect parts of responses are marked in red.
Figure 2: (a) Agent-based methods bootstrap closed-source LLMs via delicately designed prompts; (b) Instruction tuning with fixed tool-use dataset; (c) Instruction tuning with steam tool-use dataset (Ours); (d) Average tool calling accuracy on VideoToolBench vs. # learned tools. Sequential training denotes training on a sequence of tasks independently, one after the other.
@@ Introduction
Despite the progress in image-based LLMs, advancing LLMs’ capacity to comprehend video data (Lin et al. 2023a; Zhang, Li, and Bing 2023; Li et al. 2023b; Maaz et al. 2023; Li, Wang, and Jia 2023; Jin et al. 2024; Munasinghe et al. 2023; Liu et al. 2024b; Lin et al. 2023c; Jin et al. 2023; Liu et al. 2023b) still remains a demanding pursuit. The overwhelming majority of video LLM methods follow the instruction-tuning paradigm. Building upon open-source LLMs (Chiang et al. 2023; Touvron et al. 2023a, b), they entail end-to-end training on the instruction-tuning dataset generated by GPT (Achiam et al. 2023; Brown et al. 2020). These methods, however, demonstrate limited ability to comprehend video actions and present responses with hallucinations (e.g., non-existent “keyboard click” actions). Besides, most of them lack the capability of any-to-any generation, e.g., video object segmentation in Figure 1 (b).
@@
To alleviate this, agent-based methods (Gao et al. 2023; Qin et al. 2023; Yang et al. 2024b; Fan et al. 2024; Wang et al. 2024b) are introduced to advocate the tool-use capabilities of LLMs. As shown in Figure 2 (a), they mostly bootstrap closed-source LLMs (Achiam et al. 2023; Brown et al. 2020) to decompose the user instructions into more manageable sub-tasks and incorporate external tools to assist. These training-free methods require delicately designed prompts and lengthy context windows for instructions (Liu et al. 2023c). Instead, recent attempts (Liu et al. 2023c; Qin et al. 2023) explore developing tool-specific instruction tuning data and elicit the tool-use capabilities of open-source LLMs (c.f. Figure 2 (b)). However, these methods typically assume a pre-set repository containing fixed tools, which is difficult to generalize into the real-world scenario where tool data arrives in a never-ending stream. In such case, the conventional sequential training, i.e., training on new data with the pre-trained weights on the previous tool data as initialization, leads to significant catastrophic forgetting. of prior tool-usage knowledge. For example in Figure 2 (d), the tool-call accuracy of sequential training rapidly decreases as the number of tools learned increases.
@@
In this paper, we propose to enhance open-source video LLMs with COntinuaL Tool usage (dubbed as COLT). By “continual”, we mean that video LLMs possess the inherent capacity to learn and automatically activate tool functionality in a successive tool-stream format. To achieve this, two critical issues arise: 1) how to build a tool-specific memory to mitigate catastrophic forgetting of past knowledge? One intuitive idea is to store a few past tool data and replay them with the new tool data for continual training. However, the requisite memory expands proportionally with the increase in the number of tools and previous data may be unavailable due to privacy constraints. Instead, we set up a tool codebook consisting of learnable prompts to store tool-specific information in a more concise manner; 2) how to exploit relevant tools based on input user instructions? For tool activation, we compute the cosine similarity between the embedded user instructions and each prompt within the tool codebook. The highest-response prompts are then selected to “instruct” the model to invoke appropriate tools.
@@
Although several video LLM benchmarks (Maaz et al. 2023; Li et al. 2023c) have been proposed, the community still lacks the benchmark for video-centric tool-use instruction tuning. To this end, we collect VideoToolBench to facilitate tool-use capabilities for open-source video LLMs. Specifically, we collect a set of video specialist models and prompt GPT (Achiam et al. 2023) to generate diverse instructions for tool calls. Then the tool execution results are used to form the instruction-following dataset. Due to our continual learning setting, the tool list can be easily extended, making our COLT a lifelong learning video LLM.
In summary, our contributions are in three-folds:
@@
  • A tool-using instruction-following dataset VideoToolBench is introduced to unlock the potential for tool usage within video LLMs.
@@
  • Experiments on both existing video LLM benchmarks and VideoToolBench have manifested the state-of-the-art performance of COLT. For example on MSRVTT-QA (Chen and Dolan 2011), our COLT boosts the previous SOTA method (Li et al. 2023c) by 8.2% on the accuracy of zero-shot video-question answering.
Related Work
@@
Video LLMs. Recent successes of LLMs (Achiam et al. 2023; Anil et al. 2023) have shed light on the burgeoning proliferation of video LLMs (Lin et al. 2023a; Maaz et al. 2023; Gao et al. 2023). The mainstream video LLMs follow the instruction-tuning paradigm. Building upon open-source LLMs (Chiang et al. 2023; Touvron et al. 2023a, b), this kind of methods adapts the pre-trained video features into LLM understandable representations via multi-layer perception projector (Lin et al. 2023a; Maaz et al. 2023; Li, Wang, and Jia 2023; Munasinghe et al. 2023; Liu et al. 2024b; Jin et al. 2023; Liu et al. 2023b), Q-former (Zhang, Li, and Bing 2023; Li et al. 2023b), or discretization tokenizer (Jin et al. 2024). These methods, however, fail to generalize to broader video understanding tasks which may require flexible input and output formats (Wu et al. 2023; Zhan et al. 2024). Several methods attempt to achieve this by employing additional functional modules (Jin et al. 2024; Munasinghe et al. 2023) (e.g., grounding (Munasinghe et al. 2023) or diffusion modules (Jin et al. 2024)), which is not flexible enough to adapt to diverse video understanding needs. To this end, another stream of agent-based methods (Gao et al. 2023; Qin et al. 2023; Yang et al. 2024b; Fan et al. 2024; Wang et al. 2024b) is proposed, where multi-modal agents mostly bootstrap closed-source LLM (Achiam et al. 2023; Brown et al. 2020) to decompose solution paths and call external tools. This kind of method relies heavily on delicately designed prompts and may fail to acquire the usage of a novel tool. Therefore, how to implement automatic invocation of related tools in video LLMs under a non-stationary tool stream remains to be solved. This paper combines the strengths of the above two methodologies to empower video LLMs with automatic and continual tool-use ability.
@@
Continual Learning. Continual learning (Wang et al. 2024a; Lee et al. 2017; McCloskey and Cohen 1989) refers to the ability to incrementally acquire, update, and accumulate knowledge throughout the model lifetime without catastrophically forgetting previously learned information. Conceptually, four varieties of methodologies are posited. Regularization-based approaches (Kirkpatrick et al. 2017; Li and Hoiem 2017; Feng, Wang, and Yuan 2022; Yang et al. 2024a) strike the balance between the old and new tasks by adding explicit regularization terms. Architecture-based approaches  (Yoon et al. 2018; Li et al. 2019; Ke, Liu, and Huang 2020) isolate model parameters for different tasks. Rehearsal-based methods (Bonicelli et al. 2022; Chen and Chang 2023; Lin et al. 2023b) typically use a memory buffer to store several training samples from previous classes, which are used to approximate and recover old data distributions. Prompt-based methods (Wang et al. 2022c; Smith et al. 2023; Wang et al. 2022b; Li et al. 2024a) usually construct task-adaptive prompts and select appropriate prompts during inference. This kind of method is rehearsal-free and thus more computationally efficient. L2P (Wang et al. 2022c) introduces the concept of prompt pool and select prompts by a query-key mechanism. To overcome the separate optimization issue of L2P (Wang et al. 2022c), CODA-Prompt (Smith et al. 2023) assembles learnable prompts with input-conditioned weights. DualPrompt (Wang et al. 2022b) and KC-Prompt (Li et al. 2024a) set up prompt pools to respectively encode task-invariant and tasks-specific knowledge.
@@
Borrowing the favorable rehearsal-free merit, our COLT inventively devises a video LLM with the continual tool-use learning capability. We differ from current prompt-based methods (Wang et al. 2022c; Smith et al. 2023; Wang et al. 2022b; Li et al. 2024a) in the following two aspects. Firstly, our focus is on the more complex task of multi-modal language generation as opposed to the basic image classification task. Secondly, our COLT employs straight-through (Van Den Oord, Vinyals et al. 2017; Bengio, Léonard, and Courville 2013) gradient estimation to mitigate the optimization challenges encountered in previous prompt-based approaches (Wang et al. 2022c; Smith et al. 2023; Wang et al. 2022b; Li et al. 2024a).
Dataset Construction
@@
Dataset Structure. Each instance in VideoToolBench contains two rounds of conversations between human and gpt. The first round includes human instructions related to the video content and the LLM responses on choosing appropriate tools; The second round contains the execution results of the corresponding tools and the final responses from LLMs. Specifically, we follow (Yao et al. 2022; Yang et al. 2023; Liu et al. 2023c) to unify the response format of GPT into three fields including thought, action and value to mimic human-like task-solving procedure.
@@
Table 1: Tool repository of the proposed VideoToolBench including single and compositional tools.
@@
single tool
Action Recognition (AR) VideoMAE (Tong et al. 2022)
Dense Video Caption (DVC) PDVC (Wang et al. 2021)
Temporal Action Localization (TAL) InternVideo (Wang et al. 2022a)
Optical Character Recognition (OCR) EasyOCR (JaidedAI 2023)
Automatic Speech Recognition (ASR) Whisper (Radford et al. 2023)
Video Relation Detection (VRD) VidVRD (Shang et al. 2017)
Video Object Segmentation (VOS) VisTR (Wang et al. 2021)
Text-to-Video Generation (T2V) T2V (Khachatryan et al. 2023)
multi.
AR + ASR VideoMAE + Whisper
@@
Dataset Construction. This dataset is constructed using GPT-3.5-turbo (Achiam et al. 2023) with self-instruction. The involved tools [^1] In this paper, we use tool to represent the general skills and specialist model to denote the specific model. incorporate both video understanding and generation tasks. The prompt and in-context learning cases are available in supplementary materials. As shown in Table 1, we collect ten tools including eight single tool and two compositional tool. The tool list can be easily extended using a similar dataset construction manner. We initially generate 5,000 instruction-following samples for each tool. We conduct both data format checks and manual verification of semantic meanings to filter out error data. After a thorough examination, the dataset is partitioned into distinct training and testing splits, adhering to a proportionality ratio of 9:1. The full statistics of the final VideoToolBench are summarized in supplementary materials.
Method
@@
Our COLT learns the continual tool usage from the stream instruction-tuning dataset $\{\mathcal{D}^{t}\}_{t=1}^{T}$, each $\mathcal{D}^{t}$ containing triplets of the visual data $\mathbf{X}_{\mathbf{v}}^{t}$, user instructions $\mathbf{X}_{\mathbf{w}}^{t}$ and the response sequence $\mathbf{X}_{\mathbf{a}}^{t}$. For clarity, we elaborate on the architecture and training strategy for one single dataset $\mathcal{D}^{t}$ and the superscripts of $\mathbf{X}_{\mathbf{v}}^{t}$, $\mathbf{X}_{\mathbf{w}}^{t}$ and $\mathbf{X}_{\mathbf{a}}^{t}$ are omitted.
@@
Figure 3: An overview of COLT. Stage 1 aligns the visual and textual modalities through the individual training of the linear projector ${f}_{\boldsymbol{\delta}}$; In stage 2 and stage 3, the prompt within tool codebook $\mathbf{P}$ is adaptively selected according to the cosine similarity with the query feature $\mathbf{H}_{\mathbf{q}}$.
Architecture

### v3

@@ Abstract
The success of Large Language Models (LLMs) has significantly propelled the research of video understanding. To leverage the strengths of specialist models (i.e., tool) for specific video tasks, recent video LLMs have focused on integrating specialist models as tool usage capabilities into their architectures. Existing methods either prompt closed-source LLMs or employ the instruction tuning paradigm for tool-use finetuning. These methods, however, assume an established repository of fixed tools and struggle to generalize to real-world environments where tool data is perpetually evolving and streaming in. To this end, we propose to enhance open-source video LLMs with COntinuaL Tool usage (termed COLT), which automatically acquires tool-use ability in a successive tool stream without suffering “catastrophic forgetting” of the past learned tools. Specifically, our COLT incorporates a learnable tool codebook as a tool-specific memory system. Then, relevant tools are dynamically selected based on the similarity between user instructions and tool features within the codebook. To unleash the tool usage potential of video LLMs, we collect a video-centric tool-use instruction tuning dataset VideoTool. Extensive experiments on both previous video LLM benchmarks and the tool-use-specific VideoTool test split demonstrate the state-of-the-art performance of our proposed COLT.
1 Introduction
Large Language Models (LLMs), e.g., GPT (achiam2023gpt; brown2020language), PaLM (anil2023palm), and LLaMA (touvron2023llama; touvron2023llama2), have exhibited remarkable success in understanding user instructions, aligning with human intents, and generating trustworthy responses. Trained on large-scale corpora and equipped with billions of parameters, these models demonstrate impressive generalization abilities across a wide range of natural language tasks such as reasoning, summarization, and question answering. They have fundamentally reshaped the paradigm of natural language understanding by enabling unified instruction-following interfaces that support various downstream applications, from conversational agents to coding assistants. Recently, the research focus has been gradually shifting from pure text-based LLMs to multi-modal LLMs such as BLIP-2 (li2023blip), Flamingo (alayrac2022flamingo), and MiniGPT-4 (zhu2023minigpt), which augment language models with the capability to perceive and reason over multiple modalities such as vision, audio, and video. By integrating visual encoders with pretrained LLMs, these models can interpret images or videos and produce context-aware, grounded responses, thus bridging perception and language understanding. Such a shift extends the role of LLMs from passive text generators to more general perceptual agents, which has aroused significant research interest within the academic community and beyond.
@@ 1 Introduction
Despite the progress in image-based LLMs, advancing LLMs’ capacity to comprehend video data remains a demanding pursuit. Recent representative efforts such as Video-LLaMA (zhang2023video) and VideoChat (li2023videochat) illustrate the potential of video LLMs, yet the field is still in its early stage. The overwhelming majority of video LLM methods follow the instruction-tuning paradigm. Building upon open-source LLMs such as Vicuna (vicuna2023) and LLaMA (touvron2023llama; touvron2023llama2), they entail end-to-end training on instruction-tuning datasets generated by GPT models (achiam2023gpt; brown2020language). These methods, however, demonstrate limited ability to comprehend video actions and present responses with hallucinations (e.g., non-existent “keyboard click” actions). Besides, most of them lack the capability of any-to-any generation, e.g., video object segmentation in Figure 1 (b).
Figure 1: Our proposed COLT continually learns to invoke tools from tool-stream data without catastrophic forgetting. Benefiting from tool usage, COLT (a) is adept at dynamic content understanding and (b) supports flexible generation compared to existing methods lin2023video; li2023mvbench. The incorrect parts of responses are marked in red.
Figure 2: (a) Agent-based methods bootstrap closed-source LLMs via delicately designed prompts; (b) Instruction tuning with fixed tool-use dataset; (c) Instruction tuning with stream tool-use dataset (Ours); (d) Average tool calling accuracy on VideoTool vs. learned tools. Sequential training denotes training on a sequence of tasks independently.
@@
To alleviate this, agent-based methods such as AssistGPT (gao2023assistgpt), ToolLLM (qin2023toolllm), DoraemonGPT (yang2024doraemongpt), VideoAgent (fan2024videoagent), and VideoAgent-2 (wang2024videoagent) are introduced to advocate the tool-use capabilities of LLMs. More recently, VITAL (zhang2025thinkingvideosmultimodaltoolaugmented) explores reinforcing tool-augmented video reasoning via multimodal reinforcement learning on long videos. As shown in Figure 2 (a), they mostly bootstrap closed-source LLMs (achiam2023gpt; brown2020language) to decompose the user instructions into more manageable sub-tasks and incorporate external tools to assist. These training-free methods require delicately designed prompts and lengthy context windows for instructions (liu2023llava). Instead, recent attempts like LLaVA-plus (liu2023llava) and ToolLLM (qin2023toolllm) explore developing tool-specific instruction tuning data and elicit the tool-use capabilities of open-source LLMs (c.f. Figure 2 (b)). However, these methods typically assume a pre-set repository containing fixed tools, which is difficult to generalize into the real-world scenario where tool data arrives in a never-ending stream. In such cases, the conventional sequential training, i.e., training on new data with the pre-trained weights on the previous tool data as initialization, leads to significant catastrophic forgetting of prior tool-usage knowledge. For example in Figure 2 (d), the tool-call accuracy of sequential training rapidly decreases as the number of tools learned increases.
@@
In this paper, we propose to enhance open-source video LLMs with COntinuaL Tool usage (dubbed as COLT). By “continual”, we mean that video LLMs possess the inherent capacity to learn and automatically activate tool functionality in a successive tool-stream format. To achieve this, two critical issues arise: 1) how to build a tool-specific memory to mitigate catastrophic forgetting of past knowledge? One intuitive idea is to store a few past tool data and replay them with the new tool data for continual training. However, the requisite memory expands proportionally with the increase in the number of tools, and previous data may be unavailable due to privacy constraints. Instead, we set up a tool codebook consisting of learnable prompts to store tool-specific information in a more concise manner; 2) how to exploit relevant tools based on input user instructions? For tool activation, we compute the cosine similarity between the embedded user instructions and each prompt within the tool codebook. The highest-response prompts are then selected to “instruct” the model to invoke appropriate tools.
@@
Although several video LLM benchmarks (maaz2023video; li2023mvbench) have been proposed, the community still lacks a benchmark for video-centric tool-use instruction tuning. To this end, we collect VideoTool to facilitate tool-use capabilities for open-source video LLMs. Specifically, we collect a set of video specialist models and prompt GPT (achiam2023gpt) to generate diverse instructions for tool calls. Then the tool execution results are used to form the instruction-following dataset. Under our continual learning setup, COLT can incrementally incorporate newly added tools from a expanding repository, while mitigating forgetting of previously learned ones.
In summary, our contributions are threefold:
@@
  • A tool-using instruction-following dataset VideoTool is introduced to unlock the potential for tool usage within video LLMs.
@@
  • Experiments on both existing video LLM benchmarks and the VideoTool test split have manifested the state-of-the-art performance of COLT. For example, on MSRVTT-QA (chen2011collecting), our COLT boosts the previous SOTA method (li2023mvbench) by 8.2% on the accuracy of zero-shot video-question answering.
2 Related Work
@@
Video LLMs. Recent successes of LLMs (achiam2023gpt; anil2023palm) have shed light on the burgeoning proliferation of video LLMs (lin2023video; maaz2023video; gao2023assistgpt). The mainstream video LLMs follow the instruction-tuning paradigm. Building upon open-source LLMs (vicuna2023; touvron2023llama; touvron2023llama2), this kind of method adapts the pre-trained video features into LLM understandable representations via multi-layer perception projector (lin2023video; maaz2023video; li2023llama; munasinghe2023pg; liu2024st; jin2023chat; liu2023one), Q-former (zhang2023video; li2023videochat), or discretization tokenizer (jin2024video). These methods, however, fail to generalize to broader video understanding tasks, which may require flexible input and output formats (wu2023next; zhan2024anygpt). Several methods attempt to achieve this by employing additional functional modules (jin2024video; munasinghe2023pg) (e.g., grounding (munasinghe2023pg) or diffusion modules (jin2024video)), which are not flexible enough to adapt to diverse video understanding needs. To this end, another stream of agent-based methods (qin2023toolllm; yang2024doraemongpt; fan2024videoagent; wang2024videoagent) is proposed, where multi-modal agents mostly bootstrap closed-source LLM (achiam2023gpt; brown2020language) to decompose solution paths and call external tools. This kind of method relies heavily on delicately designed prompts and may fail to acquire the usage of a novel tool. Therefore, how to implement automatic invocation of related tools in video LLMs under a non-stationary tool stream remains to be solved. This paper combines the strengths of the above two methodologies to empower video LLMs with automatic and continual tool-use ability.
@@
Continual Learning. Continual learning (wang2024comprehensive; lee2017overcoming; mccloskey1989catastrophic) refers to the ability to incrementally acquire, update, and accumulate knowledge throughout the model lifetime without catastrophically forgetting previously learned information. Conceptually, four varieties of methodologies are posited. Regularization-based approaches (kirkpatrick2017overcoming; li2017learning; feng2022overcoming; yang2024CLL-CLIP) strike the balance between the old and new tasks by adding explicit regularization terms. Architecture-based approaches  (yoon2018DEN; li2019Learn2Grow; ke2020CAT) isolate model parameters for different tasks. Rehearsal-based methods (bonicelli2022effectiveness; chen2023dynamic; lin2023pcr) typically use a memory buffer to store several training samples from previous classes, which are used to approximate and recover old data distributions. Prompt-based methods (wang2022learning; smith2023coda; wang2022dualprompt; li2024kc) usually construct task-adaptive prompts and select appropriate prompts during inference. This kind of method is rehearsal-free and thus more computationally efficient. L2P (wang2022learning) introduces the concept of a prompt pool and selects prompts by a query-key mechanism. To overcome the separate optimization issue of L2P (wang2022learning), CODA-Prompt (smith2023coda) assembles learnable prompts with input-conditioned weights. DualPrompt (wang2022dualprompt) and KC-Prompt (li2024kc) set up prompt pools to respectively encode task-invariant and task-specific knowledge.
@@
Borrowing the favorable rehearsal-free merit, our COLT inventively devises a video LLM with the continual tool-use learning capability. We differ from current prompt-based methods (wang2022learning; smith2023coda; wang2022dualprompt; li2024kc) in the following two aspects. Firstly, our focus is on the more complex task of multi-modal language generation as opposed to the basic image classification task. Secondly, our COLT employs straight-through (van2017neural; bengio2013estimating) gradient estimation to mitigate the optimization challenges encountered in previous prompt-based approaches (wang2022learning; smith2023coda; wang2022dualprompt; li2024kc).
3 Dataset Construction
@@
Dataset Structure. Each instance in VideoTool contains two rounds of conversations between human and gpt. The first round includes human instructions related to the video content and the LLM responses on choosing appropriate tools; The second round contains the execution results of the corresponding tools and the final responses from LLMs. Specifically, we follow (yao2022react; yang2023mm; liu2023llava) to unify the response format of GPT into three fields, including thought, action , and value to mimic a human-like task-solving procedure.
@@
Table 1: Tool repository of the proposed VideoTool including single and compositional tools.
@@
single tool Action Recognition (AR) VideoMAE tong2022videomae
Dense Video Caption (DVC) PDVC wang2021end
Temporal Action Localization (TAL) InternVideo wang2022internvideo
Optical Character Recognition (OCR) EasyOCR (EasyOCR)
Automatic Speech Recognition (ASR) Whisper radford2023robust
Video Relation Detection (VRD) VidVRD shang2017video
Video Object Segmentation (VOS) VisTR wang2021end
Text-to-Video Generation (T2V) T2V khachatryan2023text2video
multi. AR + ASR VideoMAE + Whisper
@@
Dataset Construction. This dataset is constructed using GPT-3.5-turbo (achiam2023gpt) with self-instruction. The involved tools [^1] In this paper, we use tool to represent the general skills and specialist model to denote the specific model. incorporate both video understanding and generation tasks. The prompt and in-context learning cases are available in supplementary materials. As shown in Table 1, we collect ten tools including eight single tool and two compositional tool. The tool list can be easily extended using a similar dataset construction manner. We initially generate 5,000 instruction-following samples for each tool. We conduct both data format checks and manual verification of semantic meanings to filter out error data. After a thorough examination, the dataset is partitioned into distinct training and testing splits, adhering to a proportionality ratio of 9:1. The full statistics of the final VideoTool are summarized in supplementary materials.
4 Method
@@
Our COLT learns the continual tool usage from the stream instruction-tuning dataset $\{\mathcal{D}^{t}\}_{t=1}^{T}$, each $\mathcal{D}^{t}$ containing triplets of the visual data $\mathbf{X}_{\mathbf{v}}^{t}$, user instructions $\mathbf{X}_{\mathbf{w}}^{t}$, and the response sequence $\mathbf{X}_{\mathbf{a}}^{t}$. For clarity, we elaborate on the architecture and training strategy for one single dataset $\mathcal{D}^{t}$ and the superscripts of $\mathbf{X}_{\mathbf{v}}^{t}$, $\mathbf{X}_{\mathbf{w}}^{t}$, and $\mathbf{X}_{\mathbf{a}}^{t}$ are omitted.
@@
Figure 3: An overview of COLT. Stage 1 aligns the visual and textual modalities through the individual training of the linear projector ${f}_{\bm{\delta}}$; In stage 2 and stage 3, the prompt within tool codebook $\mathbf{P}$ is adaptively selected according to the cosine similarity with the query feature $\mathbf{H}_{\mathbf{q}}$.
4.1 Architecture